OpenStudy (yisrael012332):

Evaluate the following expression using the values given: Find 4x - y - 2z if x = -2, y = 3, and z = -2 HELP HELP HELP

OpenStudy (yisrael012332):

@ganeshie8

OpenStudy (yisrael012332):

@Hero

OpenStudy (yisrael012332):

@HWBUSTER00

OpenStudy (yisrael012332):

@dan815

OpenStudy (yisrael012332):

@dan815

OpenStudy (yisrael012332):

@princeharryyy

OpenStudy (yisrael012332):

@PrinceKel

OpenStudy (yisrael012332):

@Michele_Laino

OpenStudy (yisrael012332):

@Nnesha

OpenStudy (anonymous):

Just substitute the numbers given to you into the equation.

OpenStudy (anonymous):

\[4(-2)-(3)-2(-2)\]

OpenStudy (yisrael012332):

TY IS -7

OpenStudy (yisrael012332):

ITs##

OpenStudy (yisrael012332):

can you help with another

OpenStudy (yisrael012332):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1 over 3π(x - 7)2 (x - 3) What does the factor π (x - 7)2 represent?

OpenStudy (yisrael012332):

u have no clue'

OpenStudy (yisrael012332):

i

OpenStudy (danjs):

Radius is 4 less than the height... write that as it reads, 'is' means equals r = h - 4

OpenStudy (danjs):

Height is (x-3) h = x - 3

OpenStudy (yisrael012332):

The total surface area of the tank The area of the curved sides of the tank The area of the circular base of the tank The area of the base of about three such tanks

OpenStudy (yisrael012332):

so would it be the 3ed one

OpenStudy (danjs):

The volume of a Cone is \[V = \frac{ 1 }{ 3 }\pi*r^2*h\]

OpenStudy (princeharryyy):

insert in the values and solve

OpenStudy (danjs):

The radius is measured from the wide base of the cone

OpenStudy (yisrael012332):

oh ok so d

OpenStudy (danjs):

If the radius is height minus 4 and the height is x - 3 r = h - 4 h = x - 3 sub in h for the radius r = (x-3) - 4 = x - 7

OpenStudy (danjs):

They want to know what pi*(x-7)^2 represents... in other words now that you have r = x - 7 the question is , what is pi*r^2

OpenStudy (yisrael012332):

you lost me

OpenStudy (danjs):

ok, the givens in the problem are, r = h -4 (radius is 4 less than height) h = x - 3 (height is x - 3

OpenStudy (danjs):

you get that part?

OpenStudy (yisrael012332):

yea

OpenStudy (danjs):

To combine those two, you can take the h = x - 3, and insert that into the radius r = h - 4 = (x-3) - 4 r = x - 7 so the radius is in terms of X now, not the height

OpenStudy (danjs):

good?

OpenStudy (yisrael012332):

no

OpenStudy (danjs):

The radius is given in terms of the height of the cone... the height is given in terms of a measurememt X... is is one of those compound functions like.. r [ h(x)] You use the height value in the radius , to get r(x), the radius in terms of the measure X instead of height h

OpenStudy (yisrael012332):

i understand but What does the factor π (x - 7)2 represent

OpenStudy (yisrael012332):

i understand how you solved it but how doo i answer the question

OpenStudy (danjs):

The volume of a cone is \[V = \frac{ 1 }{ 3 }\pi*r^2*h\]

OpenStudy (danjs):

recall, the area of a circle is A = pi*r^2 So the Volume of the cone is, 1/3 * ( area of circle base) * (height)

OpenStudy (yisrael012332):

yea ok

OpenStudy (danjs):

pi times the radius squared is the area of the base in this instance

OpenStudy (yisrael012332):

so its c

OpenStudy (yisrael012332):

like i said

OpenStudy (danjs):

\[V = \frac{ 1 }{ 3 }\pi*r^2*h = (area ~base)*\frac{ height }{ 3 }\]

OpenStudy (danjs):

i didnt look at the answers

OpenStudy (yisrael012332):

The total surface area of the tank The area of the curved sides of the tank The area of the circular base of the tank The area of the base of about three such tanks

OpenStudy (yisrael012332):

i understand the problem no but i wanted to clarify

OpenStudy (danjs):

a, right on, yeah C

OpenStudy (yisrael012332):

tytytytytyty so much